It takes Dan 20 minutes to mow the lawn; it takes his sister,Sue, 15 minutes to mow the same lawn. If they work together, how long will it take to mow the lawn?
Answer by jorel1380(2518)

1/20+1/15=1/t
3t+4t=60
7t=60
t=60/7 minutes..